Janes Company provided the following information on intangible assets:

  1. A patent was purchased from the Lou Company for $1,550,000 on January 1, 2019. Janes estimated the remaining useful life of the patent to be 10 years. The patent was carried on Lou’s accounting records at a net book value of $520,000 when Lou sold it to Janes.
  2. During 2021, a franchise was purchased from the Rink Company for $670,000. The contractual life of the franchise is 10 years and Janes records a full year of amortization in the year of purchase.
  3. Janes incurred research and development costs in 2021 as follows:
Materials and supplies $ 157,000
Personnel 197,000
Indirect costs 77,000
Total $ 431,000
  1. Effective January 1, 2021, based on new events that have occurred, Janes estimates that the remaining life of the patent purchased from Lou is only five more years.


Required:
1. Prepare the entries necessary for years 2019 through 2021 to reflect the above information.

  • 1

    Record the purchase of a patent.

  • 2

    Record amortization on the patent.

  • 3

    Record amortization on the patent.

  • 4

    Record the purchase of a franchise.

  • 5

    Record amortization of franchise.

  • 6

    Record research and development expenses.

  • 7

    Record amortization on the patent after change in useful life.

2. Prepare a schedule showing the intangible asset section of Janes’s December 31, 2021, balance sheet.

Prepare a schedule showing the intangible asset section of Janes’s December 31, 2021, balance sheet.

Partial Balance Sheet
December 31, 2021
Intangible assets:
  
Total intangibles
Answer #1

Solution

Janes Company

Journal Entries:

Date

Account Titles and Explanation

Debit

Credit

Jan 1, 2019

Patent

$1,550,000

Cash

$1,550,000

(To record the purchase of patent)

Dec 31, 2019

Amortization Expense - Patent

$155,000

Accumulated Amortization - Patent

$155,000

(To record amortization on patent; 1,550,000/10 years = $155,000)

Dec 31, 2020

Amortization Expense - Patent

$155,000

Accumulated Amortization - Patent

$155,000

(To record amortization on patent; 1,550,000/10 years = $155,000)

Jan 1, 2021

Franchise

$670,000

Cash

$670,000

(To record purchase of franchise)

Jan 1, 2021

Research and Development Expenses

$431,000

Cash

$431,000

(To record research and development expenses)

Dec 31, 2021

Amortization Expense - Franchise

$67,000

Accumulated Amortization - Franchise

$67,000

(To record amortization expense on franchise; 670,000/10 = $67,000)

Dec 31, 2021

Amortization Expense - Patent

$248,000

Accumulated Amortization - Patent

$248,000

(To record amortization expense on patent)

Note – Computation of patent amortization on December 31, 2021:

Book value = cost – accumulated amortization

= $1,550,000 – (155,000 + 155,000) = $1,240,000

Remaining useful life of patents = 5 years

Annual amortization = $1,240,000/5 = $248,000

  1. Schedule showing the intangible assets section of Janes’ December 31, 2021 balance sheet:

Janes Company

Balance Sheet (Partial) at December 31, 2021

Assets:

Intangible Assets

Patents

$1,550,000

Accumulated Amortization

$558,000

$992,000

Franchise

$670,000

Accumulated Amortization

$67,000

$603,000

Total Intangible Assets

$1,595,000